Hydrogen gas was passed through water and the total total pressure was 0.986 bar at 25 degrees C.
Calculate the partial pressure of hydrogen in Torr. Enter a number with correct significant figures and DO NOT include units as they are already indicated.
Calculate the mole percent of hydrogen. Enter the value as a percent with the correct significant figures. DO NOT include units as they are already indicated.
Calculate the root-mean-squared speed of hydrogen gas at 25 degrees C. Enter the value as a percent with the correct significant figures. DO NOT include units as they are already indicated.
1
Expert's answer
2020-11-23T07:33:12-0500

0.986 bar = 739.56torr

Total pressure = partial pressure of the dry hydrogen gas + partial pressure of water vapor

= 739.56 + 23.8 = 763.36 torr

% = 739.56/ 763.36 × 100 = 76.88


RMS= square root(3RT/M) = square root ( 3 ×8.314 ×298 / 2×10-3) = 1927.79